Hello the problem I have is in the multiplication of polynomials in the book it is: An athletic field has dimensions 30 yd by 100 yd. An end zone that is w yards wide borders each end of the field. express the total area of the field and the end zones in terms of the variable w. Thank you and I appreciate any help.
Answer by stanbon(75887)

Hello the problem I have is in the multiplication of polynomials in the book it is: An athletic field has dimensions 30 yd by 100 yd. An end zone that is w yards wide borders each end of the field. express the total area of the field and the end zones in terms of the variable w.
-----
Draw the picture::
The length of the field including the end zones is 100 +2w
The width of the field is 30
---------------
Area = width*length = 30(100+2w) = 3000 + 60w
Ans: A = 3000 + 60w
-------------------------
Cheers,
Stan H.
------------------